OpenStudy (anonymous):

(5/2x+6)-2=(1-8x/4x) x=?

OpenStudy (anonymous):

(5/2x+6)-2 = 1 - 2 (5/2x+6)-2 = -1 (5/2x+6) = 1 5/2x+6 = 1 5/2x = -5 (for explanation why the next step is allowed, try rewrting the equation when the variables were 6/3 = 2) 5/-5 = 2x from here on it shouldnt be a problem

OpenStudy (anonymous):

-1/2?